ToxicaDennisIce:

Which expression is equivalent to 6 (a minus 3)?

ToxicaDennisIce:

jhonyy9:

so there is 6(a-3) how you distribute parentheses ?

ToxicaDennisIce:

um

ToxicaDennisIce:

help

jimthompson5910:

Here's a similar example 3(x-10) We multiply the outer 3 by each term inside (the x and -10) 3 times x = 3x = 3(x) 3 times -10 = 3(-10) = -3(10) This means 3(x-10) = 3(x) - 3(10) So on paper you could write it like this 

jimthompson5910:

Then distribute the 3 through like so